OpenStudy (anonymous):

It takes Tom 15 minutes to wash a car.It takes only 10 minutes with jerry helping.how long should it take jerry to wash the car by himself? A. 6 minutes b. 12.5 minutes c. none of these choices are correct. d. 25 minutes e. 5 minutes

OpenStudy (anonymous):

would take 30 minutes

OpenStudy (anonymous):

c

OpenStudy (saifoo.khan):

How'd you got it?

OpenStudy (saifoo.khan):

1/15 + 1/x = 1/10

OpenStudy (anonymous):

1/15 + 1/x = 1/10

OpenStudy (mertsj):

Tom can wash 1/15 of the car in 1 minute Together they wash 1/10 of the car in 1 minutes. The part Tom does + the part Jerry does = the part they do together. So \[\frac{1}{15}+\frac{1}{x}=\frac{1}{10}\]

OpenStudy (anonymous):

can you solve that iheart?

OpenStudy (anonymous):

no sir i can not

OpenStudy (mertsj):

Multiply each term by the common denominator which is 30x

OpenStudy (anonymous):

ok - first find LCM of 10,15 and x this is 30x multiply each term by 30x gives 30x 30x 30x ---- + --- = --- 15 x 10 dividing each term 2x + 30 = 3x 30 = 3x - 2x = x so x = 30 none of the choices are correct - as kumar said the answer is c

OpenStudy (anonymous):

thank very kind

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

yw